# AraFinNews: Arabic Financial Summarisation with Domain-Adapted LLMs

### v1

@@ Abstract
This paper investigates the impact of domain specificity on abstractive summarisation of Arabic financial texts using large language models (LLMs). We introduce AraFinNews, the largest publicly available Arabic financial news dataset to date, comprising 212,500 article–headline pairs spanning nearly a decade of reporting from October 2015 to July 2025. Designed as the Arabic equivalent of major English summarisation corpora such as CNN/DailyMail, AraFinNews provides a robust benchmark for evaluating domain-specific language understanding and generation in financial contexts. Using this resource, we evaluate transformer-based models—including mT5, AraT5, and the domain-adapted FinAraT5—to examine how financial-domain pretraining influences factual accuracy, numerical reliability, and stylistic alignment with professional reporting. Experimental results show that domain-adapted models generate more faithful and coherent summaries, particularly in handling quantitative and entity-centric information. The findings highlight the importance of domain-specific adaptation for improving factual consistency and narrative fluency in Arabic financial summarisation. The dataset is freely available for non-commercial research athttps://github.com/ArabicNLP-uk/AraFinNews
@@ I Introduction
Arabic, spoken by over 300 million people across more than twenty countries [1], holds central importance in religious, political, and cultural communication throughout the Arab world. Despite its global prominence, Arabic remains underrepresented in Natural Language Processing (NLP). While initiatives such as ACL SIGARAB [^1] https://www.sigarab.org/ have improved visibility, progress in building robust, domain-specific Arabic language models and datasets continues to lag behind high-resource languages such as English and Chinese [2]. This gap stems not only from data scarcity but also from the language’s structural complexity: Arabic is diglossic, encompassing Classical Arabic, Modern Standard Arabic (MSA), and diverse dialects that differ lexically, morphologically, and syntactically. Moreover, its right-to-left script, absence of capitalisation, and omission of diacritics—short vowel markers that disambiguate homographs—pose challenges for tokenisation, parsing, and semantic interpretation [3].
@@
Recent advances in Arabic NLP through models such as AraBERT [4], ARBERT, MARBERT [5], and generative models like AraBART and AraT5 [6, 7] have substantially improved general text understanding. Yet, these models are trained on broad web or news data, lacking exposure to specialised financial or technical language. In contrast, English NLP has benefited from domain-adapted architectures such as FinBERT [8] and BioBERT [9], which deliver notable performance gains in financial and biomedical applications. The absence of comparable domain adaptation for Arabic highlights the need for targeted pretraining strategies to capture the linguistic and conceptual nuances of financial discourse.
@@
The Arab financial sector has rapidly expanded over the past two decades, with markets such as Tadawul [^2] https://www.saudiexchange.sa, the Dubai Financial Market (DFM) [^3] https://www.dfm.ae, and the Abu Dhabi Securities Exchange (ADX) [^4] https://www.adx.ae now ranking among the world’s most active [10]. This growth has generated a surge of Arabic financial narratives—ranging from corporate disclosures and regulatory filings to market analyses and news reports—typically long, information-dense, and stylistically formal, making them ideal for automatic summarisation [11]. Yet most Arabic summarisation systems focus on general news, leaving financial text underexplored despite its demand for factual accuracy, numerical precision, and stylistic clarity [12, 13, 14].
@@
To address this gap, we introduce AraFinNews—the largest publicly available Arabic financial news dataset, comprising 212,500 article–headline pairs from reputable financial media sources. Designed as the Arabic equivalent of major English summarisation benchmarks such as CNN/DailyMail, AraFinNews provides a rich domain-specific resource for abstractive summarisation and related financial NLP tasks. It is freely available for research at https://github.com/ArabicNLP-UK/AraFinNews. Using this dataset, we evaluate transformer-based architectures—including FinAraT5, AraT5, and mT5—on Arabic financial summarisation, examining how domain-adapted large language models enhance fluency, factual consistency, and contextual relevance. This work contributes to advancing domain-specific Arabic NLP, demonstrating the benefits and challenges of adapting large language models to one of the world’s most linguistically complex and economically dynamic languages.
@@ II Related Work
Text summarisation has long been a key task in Natural Language Processing (NLP), particularly in journalism and information retrieval. Early systems were largely extractive, relying on sentence ranking and statistical heuristics such as term frequency, cue words, and sentence position [15, 16]. While these methods captured surface salience, they lacked true abstraction and often produced redundant or disjointed summaries.
@@
The emergence of neural encoder–decoder architectures marked a significant shift. Models based on recurrent neural networks (RNNs) with attention mechanisms introduced the ability to generate novel, coherent text rather than copy input segments verbatim. One of the first neural systems capable of producing fluent news headlines using LSTM-based architectures was demonstrated in [17]. Later developments integrated coverage mechanisms and reinforcement learning to address factual consistency and redundancy [18, 19].
@@
Transformer-based models, notably BART and T5, further revolutionised abstractive summarisation. By treating all NLP tasks as text-to-text transformations, these architectures achieved strong generalisation across domains and languages [20]. Their multilingual variants, such as mT5 [21], extended these capabilities to low-resource languages through cross-lingual transfer. While effective in general domains, these models often struggle in specialised contexts—such as finance—where domain-specific terminology, numerical reasoning, and formal discourse require more focused adaptation.
@@
In recent years, Arabic NLP has seen rapid progress due to the introduction of large-scale pretrained language models. Encoder-based architectures like AraBERT [4], ARBERT and MARBERT[5], have become benchmarks for classification and sentiment analysis tasks, addressing key challenges such as morphology and dialectal variation. Generative architectures including AraBART and AraT5 [6, 7], and more recent efforts such as Jasmine [22], have extended these capabilities to summarisation and text generation.
@@
A growing body of work now focuses on Arabic abstractive summarisation. [23] demonstrated the effectiveness of fine-tuning multilingual transformers for Arabic news summarisation, while [24] proposed transformer-based architectures that improved both coherence and fluency over traditional models. Other work has explored integrating linguistic knowledge or auxiliary tasks: for instance, [25] combined named entity recognition with summarisation to better preserve factual content, and [26] developed a hybrid RNN–Transformer model to improve grammatical consistency. These advances collectively demonstrate the viability of generative models for Arabic, though the majority of research remains limited to general news rather than specialised professional domains.
@@
Considerable advances have been made in Arabic language modelling with the introduction of pretrained systems such as AraBERT [4], ARBERT [5], MARBERT, and generative variants like AraBART and AraT5 [6, 7]. However, these remain largely general-purpose, drawing on broad web and news corpora rather than domain-specific text. In English, models like FinBERT [8] and BioBERT [9] have established the effectiveness of domain-oriented pretraining, yet equivalent efforts for Arabic are still limited. Existing Arabic models show strong performance in generic understanding tasks but struggle with the precision and structure required for financial narrative summarisation.
@@ II-A Datasets for Arabic Summarisation and Financial Corpora
Research in Arabic summarisation has been supported by a range of corpora covering both extractive and abstractive paradigms. Early efforts such as Kalimat [27], a 20,000-article Omani newspaper corpus, and the Essex Arabic Summaries Corpus (EASC) [2], which includes 153 news articles and 765 human-written summaries, provided the foundation for extractive summarisation. More recent multilingual datasets have enabled progress in neural and cross-lingual approaches. The Arabic Gigaword corpus [28], although large and widely used, lacks aligned document–summary pairs and remains commercially restricted, limiting its direct use for supervised learning. By contrast, XLSum [29] and WikiLingua [30] introduced openly available abstractive benchmarks with approximately 40,000 and 21,000 Arabic examples respectively—though their content and style differ from journalistic and financial text. Additional corpora such as SumArabic, which offers over 84,000 abstractive summaries, and HASD, containing 43,000 hybrid samples, have expanded the scale of available data but still show inconsistencies in length and domain coverage.
@@
Early Arabic summarisation research applied extractive techniques including TF–IDF ranking, lexical scoring, and graph-based centrality measures [31, 32, 33, 34]. More recently, multilingual models such as mBART and mT5 have been fine-tuned for Arabic summarisation [35], achieving improvements in lexical diversity and fluency but still facing limitations in factual precision and domain adaptation.
Unlike English, Arabic still lacks large, well-curated corpora for specialised domains. Existing resources such as BORSAH [36] and the Arabic Business and Management Corpus (ABMC) [^5] https://elhaj.uk/corpora.html provide valuable text collections in finance and economics, but they were primarily designed for sentiment analysis or classification rather than summarisation. As a result, progress in Arabic abstractive summarisation—particularly within financial contexts—has been constrained by the absence of domain-focused, large-scale, high-quality datasets.
@@ III AraFinNews: The Arabic Financial News Dataset
To address the lack of large, domain-specific resources for Arabic summarisation, we introduce AraFinNews—the largest publicly available Arabic financial news dataset to date. The corpus comprises 212,500 article–headline pairs collected from Argaam.com, one of the leading financial media platforms in the Gulf region. Each headline functions as an abstractive summary of its corresponding article, mirroring the concise and information-dense style of professional Arabic financial journalism. The dataset spans corporate disclosures, market analyses, and macroeconomic reports, offering a robust benchmark for abstractive summarisation and related NLP tasks such as named entity recognition, event extraction, and sentiment analysis.
@@
All content was collected in accordance with Argaam’s publicly accessible content policy and complies with ethical data usage standards. Argaam’s robots.txt file does not prohibit web scraping, and all retrieved material was publicly available without subscription or authentication. Data collection was restricted to factual news articles written in Modern Standard Arabic, excluding opinion pieces and proprietary editorial content, consistent with established academic practice for web-sourced corpora. The released dataset includes comprehensive metadata fields—article ID, title, date, full text, and source URL—to facilitate transparent and reproducible research.
@@ III-H Dataset Comparison and Abstractiveness Analysis
To better contextualise the characteristics of our dataset, we compare the AraFinNews corpus with well-established English summarisation datasets, including CNN, DailyMail, and the Arabic financial summarisation benchmark FinAraSum [37]. Table II summarises the basic statistics, while Table III reports abstractiveness and baseline performance using the LEAD-1 metric.
AraFinNews is comparable in nature to FinAraSum, both being derived from financial newswire content where the article body and headline reflect factual, domain-specific reporting. However, AraFinNews contains over 210K articles—substantially larger than FinAraSum (44.9K).
@@
Overall, these statistics confirm that AraFinNews shares the core properties of FinAraSum—clean, factual content and high abstractiveness—while offering broader coverage and scale. Its characteristics make it a valuable resource for benchmarking Arabic headline generation and for evaluating domain-specific language models such as FinAraT5.
Dataset Train / Val / Test Avg. Doc. Len. Avg. Sum. Len. Vocab (Docs) Vocab (Sums) CNN 90,266 / 1,220 / 1,093 760.5 45.7 34K 89K DailyMail 197,461 / 12,148 / 10,397 653.3 54.6 564K 180K FinAraSum 44,900 / 2,800 / 2,800 238.3 9.0 493K 79K AraFinNews (ours) 170,000 / 21,250 / 21,250 188 11.22 528K 83K
TABLE II: Comparison of dataset statistics between AraFinNews and existing summarisation datasets. Train/validation/test sizes are given in number of documents. Average document and summary lengths are in words. Vocabulary sizes (unique tokens) are shown in thousands. Figures for English datasets are based on standard splits reported in prior work [38].
@@ IV Financial Large Language Models for Arabic Summarisation
Recent advancements in LLMs have transformed financial text processing, enabling sophisticated applications such as document summarisation, sentiment analysis, and report generation. Prominent models such as GPT [39], PaLM [40], and FLAN-T5 [41] have achieved state-of-the-art performance in English and multilingual settings. However, their effectiveness in Arabic remains limited by the scarcity of large, high-quality domain-specific corpora and by the inherent linguistic challenges of the language—particularly its morphological richness, diglossic variation, and orthographic ambiguity.
To address this gap, Arabic-adapted transformer models such as AraBERT [42] and AraT5 [7] have extended large-scale pretraining to Modern Standard Arabic. More recently, domain-specialised variants such as FinAraT5 [37] have been developed to capture the unique stylistic and numerical features of financial discourse.
In this study, we employ FinAraT5777https://huggingface.co/drelhaj/FinAraT5 alongside multilingual and general-purpose Arabic transformers to perform abstractive summarisation on AraFinNews. These models are evaluated under a unified experimental setup designed to assess their ability to generate accurate, coherent, and factually faithful financial headlines. The following subsections outline the model selection, fine-tuning configuration, and evaluation framework adopted in this work.

### v2

@@ Abstract
This paper investigates the impact of domain specificity on abstractive summarisation of Arabic financial texts using large language models (LLMs). We introduce AraFinNews, the largest publicly available Arabic financial news dataset to date, comprising 212,500 article–headline pairs spanning nearly a decade of reporting from October 2015 to July 2025. Designed as the Arabic equivalent of major English summarisation corpora such as CNN/DailyMail, AraFinNews provides a robust benchmark for evaluating domain-specific language understanding and generation in financial contexts. Using this resource, we evaluate transformer-based models—including mT5, AraT5, and the domain-adapted FinAraT5—to examine how financial-domain pretraining influences accuracy, numerical reliability, and stylistic alignment with professional reporting. Experimental results show that domain-adapted models generate more coherent summaries, particularly in handling quantitative and entity-centric information. The findings highlight the importance of domain-specific adaptation for improving narrative fluency in Arabic financial summarisation. The dataset is freely available for non-commercial research athttps://github.com/ArabicNLP-uk/AraFinNews
@@ I Introduction
Arabic, spoken by over 300 million people across more than twenty countries [statista2024], holds central importance in religious, political, and cultural communication throughout the Arab world. Despite its global prominence, Arabic remains underrepresented in Natural Language Processing (NLP). While initiatives such as ACL SIGARAB [^1] https://www.sigarab.org/ have improved visibility, progress in building robust, domain-specific Arabic language models and datasets continues to lag behind high-resource languages such as English and Chinese [el2010using]. This gap stems not only from data scarcity but also from the language’s structural complexity: Arabic is diglossic, encompassing Classical Arabic, Modern Standard Arabic (MSA), and diverse dialects that differ lexically, morphologically, and syntactically. Moreover, its right-to-left script, absence of capitalisation, and omission of diacritics—short vowel markers that disambiguate homographs—pose challenges for tokenisation, parsing, and semantic interpretation [el-haj-rayson-2016-osman].
@@
Recent advances in Arabic NLP through models such as AraBERT [antoun_arabert_2021], ARBERT, MARBERT [abdul-mageed_arbert_2021], and generative models like AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022] have substantially improved general text understanding. Yet, these models are trained on broad web or news data, lacking exposure to specialised financial or technical language. In contrast, English NLP has benefited from domain-adapted architectures such as FinBERT [yang2020finbert] and BioBERT [lee2020biobert], which deliver notable performance gains in financial and biomedical applications. The absence of comparable domain adaptation for Arabic highlights the need for targeted pretraining strategies to capture the linguistic and conceptual nuances of financial discourse.
@@
The Arab financial sector has rapidly expanded over the past two decades, with markets such as Tadawul [^2] https://www.saudiexchange.sa, the Dubai Financial Market (DFM) [^3] https://www.dfm.ae, and the Abu Dhabi Securities Exchange (ADX) [^4] https://www.adx.ae now ranking among the world’s most active [zmandar2021joint]. This growth has generated a surge of Arabic financial narratives—ranging from corporate disclosures and regulatory filings to market analyses and news reports—typically long, information-dense, and stylistically formal, making them ideal for automatic summarisation [zmandar2021multilingual]. Yet most Arabic summarisation systems focus on general news, leaving financial text underexplored despite its demand for accuracy, numerical precision, and stylistic clarity [el2020financial, zmandar2021financial, zavitsanos2023financial].
@@
To address this gap, we introduce AraFinNews—the largest publicly available Arabic financial news dataset, comprising 212,500 article–headline pairs from reputable financial media sources. Designed as the Arabic equivalent of major English summarisation benchmarks such as CNN/DailyMail, AraFinNews provides a rich domain-specific resource for abstractive summarisation and related financial NLP tasks. It is freely available for research at https://github.com/ArabicNLP-UK/AraFinNews. Using this dataset, we evaluate transformer-based architectures—including FinAraT5, AraT5, and mT5—on Arabic financial summarisation. This work contributes to advancing domain-specific Arabic NLP, demonstrating the benefits and challenges of adapting large language models to one of the world’s most linguistically complex and economically dynamic languages.
@@ II Related Work
Text summarisation has long been a key task in Natural Language Processing (NLP), particularly in journalism and information retrieval. Early systems were largely extractive, relying on sentence ranking and statistical heuristics such as term frequency, cue words, and sentence position [filippova2009company, berger2000ocelot]. While these methods captured surface salience, they lacked true abstraction and often produced redundant or disjointed summaries.
@@
The emergence of neural encoder–decoder architectures marked a significant shift. Models based on recurrent neural networks (RNNs) with attention mechanisms introduced the ability to generate novel, coherent text rather than copy input segments verbatim. One of the first neural systems capable of producing fluent news headlines using LSTM-based architectures was demonstrated in [lopyrev2015generating]. Later developments integrated coverage mechanisms and reinforcement learning to address factual consistency and redundancy [see-etal-2017-get, cheng-etal-2016-neural].
@@
Transformer-based models, notably BART and T5, further revolutionised abstractive summarisation. By treating all NLP tasks as text-to-text transformations, these architectures achieved strong generalisation across domains and languages [raffel_exploring_2020]. Their multilingual variants, such as mT5 [xue-etal-2021-mt5], extended these capabilities to low-resource languages through cross-lingual transfer. While effective in general domains, these models often struggle in specialised contexts—such as finance—where domain-specific terminology, numerical reasoning, and formal discourse require more focused adaptation.
@@
In recent years, Arabic NLP has seen rapid progress due to the introduction of large-scale pretrained language models. Encoder-based architectures like AraBERT [antoun_arabert_2021], ARBERT and MARBERT[abdul-mageed_arbert_2021], have become benchmarks for classification and sentiment analysis tasks, addressing key challenges such as morphology and dialectal variation. Generative architectures including AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022], and more recent efforts such as Jasmine [abdul2023jasmine], have extended these capabilities to summarisation and text generation.
@@
A growing body of work now focuses on Arabic abstractive summarisation. [kahla2023fine] demonstrated the effectiveness of fine-tuning multilingual transformers for Arabic news summarisation, while [morsi2025arabic] proposed transformer-based architectures that improved both coherence and fluency over traditional models. Other work has explored integrating linguistic knowledge or auxiliary tasks: for instance, [essa2025enhanced] combined named entity recognition with summarisation to better preserve factual content, and [elsaid2025hybrid] developed a hybrid RNN–Transformer model to improve grammatical consistency. These advances collectively demonstrate the viability of generative models for Arabic, though the majority of research remains limited to general news rather than specialised professional domains.
@@
Considerable advances have been made in Arabic language modelling with the introduction of pretrained systems such as AraBERT [antoun_arabert_2021], ARBERT [abdul-mageed_arbert_2021], MARBERT, and generative variants like AraBART and AraT5 [eddine_arabart_2022, nagoudi_arat5_2022]. However, these remain largely general-purpose, drawing on broad web and news corpora rather than domain-specific text. In English, models like FinBERT [yang2020finbert] and BioBERT [lee2020biobert] have established the effectiveness of domain-oriented pretraining, yet equivalent efforts for Arabic are still limited. Existing Arabic models show strong performance in generic understanding tasks but struggle with the precision and structure required for financial narrative summarisation.
@@ II-A Datasets for Arabic Summarisation and Financial Corpora
Research in Arabic summarisation has been supported by a range of corpora covering both extractive and abstractive paradigms. Early efforts such as Kalimat [ElHaj2013KALIMATAM], a 20,000-article Omani newspaper corpus, and the Essex Arabic Summaries Corpus (EASC) [el2010using], which includes 153 news articles and 765 human-written summaries, provided the foundation for extractive summarisation. More recent multilingual datasets have enabled progress in neural and cross-lingual approaches. The Arabic Gigaword corpus [Parker2019gigawords], although large and widely used, lacks aligned document–summary pairs and remains commercially restricted, limiting its direct use for supervised learning. By contrast, XLSum [hasan-etal-2021-xl] and WikiLingua [ladhak-etal-2020-wikilingua] introduced openly available abstractive benchmarks with approximately 40,000 and 21,000 Arabic examples respectively—though their content and style differ from journalistic and financial text. Additional corpora such as SumArabic, which offers over 84,000 abstractive summaries, and HASD, containing 43,000 hybrid samples, have expanded the scale of available data but still show inconsistencies in length and domain coverage.
@@
Early Arabic summarisation research applied extractive techniques including TF–IDF ranking, lexical scoring, and graph-based centrality measures [El-Haj-2011, Douzidia2004LakhasAA, el2013using, koulali2013arabic]. More recently, multilingual models such as mBART and mT5 have been fine-tuned for Arabic summarisation [kahla-etal-2021-cross], achieving improvements in lexical diversity and fluency.
Unlike English, Arabic still lacks large, well-curated corpora for specialised domains. Existing resources such as BORSAH [ALSHAHRANI18] and the Arabic Business and Management Corpus (ABMC) [^5] https://elhaj.uk/corpora.html provide valuable text collections in finance and economics, but they were primarily designed for sentiment analysis or classification rather than summarisation. As a result, progress in Arabic abstractive summarisation—particularly within financial contexts—has been constrained by the absence of domain-focused, large-scale, high-quality datasets.
@@ III AraFinNews: The Arabic Financial News Dataset
To address the lack of large, domain-specific resources for Arabic summarisation, we introduce AraFinNews, the largest publicly available Arabic financial news dataset to date. The corpus comprises 212,500 article–headline pairs sourced through Argaam.com, a major financial media platform in the Gulf region that aggregates material from several well-established outlets, including the Financial Times, Forbes, the World Gold Council and Bloomberg. Each headline serves as an abstractive summary of its corresponding article, reflecting the concise, information-rich style characteristic of professional Arabic financial journalism. The dataset spans corporate disclosures, market analyses and macroeconomic reports, offering a solid benchmark for abstractive summarisation as well as related NLP tasks such as named entity recognition, event extraction and sentiment analysis.
@@
All content was collected in accordance with Argaam’s publicly accessible content policy and complies with ethical data usage standards. Argaam’s robots.txt file does not prohibit web scraping, and all retrieved material was publicly available without subscription or authentication. The released dataset includes comprehensive metadata fields—article ID, title, date, full text, and source URL—to facilitate transparent and reproducible research.
@@ III-H Dataset Comparison and Abstractiveness Analysis
To better contextualise the characteristics of our dataset, we compare the AraFinNews corpus with well-established English summarisation datasets, including CNN, DailyMail, and the Arabic financial summarisation benchmark FinAraSum [zmandar2023finarat5]. Table II summarises the basic statistics, while Table III reports abstractiveness and baseline performance using the LEAD-1 metric.
AraFinNews is comparable in nature to FinAraSum, both being derived from financial newswire content where the article body and headline reflect domain-specific reporting. However, AraFinNews contains over 210K articles—substantially larger than FinAraSum (44.9K).
@@
These statistics confirm that AraFinNews shares the core properties of FinAraSum abstractiveness—while offering broader coverage and scale. Its characteristics make it a valuable resource for benchmarking Arabic headline generation and for evaluating domain-specific language models such as FinAraT5.
Dataset Train / Val / Test Avg. Doc. Len. Avg. Sum. Len. Vocab (Docs) Vocab (Sums) CNN 90,266 / 1,220 / 1,093 760.5 45.7 340K 89K DailyMail 197,461 / 12,148 / 10,397 653.3 54.6 564K 180K FinAraSum 44,900 / 2,800 / 2,800 238.3 9.0 493K 79K AraFinNews (ours) 170,000 / 21,250 / 21,250 188 11.22 528K 83K
TABLE II: Comparison of dataset statistics between AraFinNews and existing summarisation datasets. Train/validation/test sizes are given in number of documents. Average document and summary lengths are in words. Vocabulary sizes (unique tokens) are shown in thousands. Figures for English datasets are based on standard splits reported in prior work [kamal-eddine-etal-2021-barthez].
@@ IV Financial Large Language Models for Arabic Summarisation
Recent advancements in LLMs have transformed financial text processing, enabling sophisticated applications such as document summarisation, sentiment analysis, and report generation. Prominent models such as GPT [brown2020language], PaLM [chowdhery2022palm], and FLAN-T5 [chung2022scaling] have achieved state-of-the-art performance in English and multilingual settings. However, their effectiveness in Arabic remains limited by the scarcity of large, high-quality domain-specific corpora and by the inherent linguistic challenges of the language—particularly its morphological richness, diglossic variation, and orthographic ambiguity.
To address this gap, Arabic-adapted transformer models such as AraBERT [antoun2020arabert] and AraT5 [nagoudi_arat5_2022] have extended large-scale pretraining to Modern Standard Arabic. More recently, domain-specialised variants such as FinAraT5 [zmandar2023finarat5] have been developed to capture the unique stylistic and numerical features of financial discourse.
In this study, we employ FinAraT5777https://huggingface.co/drelhaj/FinAraT5 alongside multilingual and general-purpose Arabic transformers to perform abstractive summarisation on AraFinNews. These models are evaluated under a unified experimental setup designed to assess their ability to generate coherent financial headlines. The following subsections outline the model selection, fine-tuning configuration, and evaluation framework adopted in this work.